OpenStudy (anonymous):

a cube has a surface area of 3750 square feet. What is its volume? Please show your work

OpenStudy (anonymous):

surface area of a cube = 6(side)^2 so 3750 = 6(side)^2 3750/6 = side^2 625 = side^2 sqrt(625) = side 25 = side ft. volume of cube = side^3 = 25^3 = 15625 cu.ft.